# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 8. Weak solutions of ISDE for the lower Dirichlet form
Throughout this proof, $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ will be denoted by the same symbol $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ and $\mathsf{X}$ will denote the unlabeled dynamics given by $\mathbf{X}$. Let $\mathsf{K}[a_{\mathsf{q}}]$ be the compact set defined by (13.16):
$\mathsf{K}[a_{\mathsf{q}}]=\bigcap_{R=1}^{\infty}\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(S_{R})\leq a_{\mathsf{q}}(R)\}.$
Let $\mathsf{S}_{\mathrm{s}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\{{x}\})\in\{0,1\}\text{ for all }{x}\in\mathbb{R}^{d}\}$. Then, according to (13.18)–(13.20), (10.3) follows from the following:
$P(\lim_{\mathsf{q}\to\infty}\kappa_{\mathsf{q}}=\infty)=1.$ (10.5)
[…]
$\{\mathbf{S}_{\mu_{\mathsf{a}}}\}$ is also clear because the $m$-labeled process $\mathsf{X}^{[m]}$ given by $\mathbf{X}$ is associated with a symmetric Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
[…]
Let $\mathbf{M}=(M^{i})_{i=1}^{\infty}$ such that $M_{t}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}$. Then, because $\mathfrak{a}=\sigma^{t}\sigma$, $\sigma\in C_{b}^{1}(\mathbb{R}^{d})$, and $\mathfrak{a}$ is uniformly elliptic, we can represent $\mathbf{B}=(B^{i})_{i=1}^{\infty}$ as a function of $\mathbf{M}$. Thus, $\mathbf{B}$ is a $\sigma[\mathbf{M}_{s};s\leq t]$-adapted process. Because $(\mathbf{X},\mathbf{B})$ is a solution of ISDE (2.41),
We can easily find $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ in (13.16) that satisfies (13.17) and (13.31). Therefore, we can use (13.30) with reference to Lemma 13.7. Thus, we have verified all assumptions of Lemma 13.8, which yields (10.5). We have thus completed the proof. ∎
Lemma 10.9.
[…]
Proof.
Let $\mathfrak{u}^{[m,0]}\!:\!(\mathbb{R}^{d})^{m}\times\mathsf{S}\!\to\!\mathsf{S}$ be the map such that $\mathfrak{u}^{[m,0]}(\mathbf{x},\mathsf{s})=\mathfrak{u}(\mathbf{x})+\mathsf{s}$.
Let $\mathsf{S}_{\mathrm{sde}}$ be the set introduced before (13.14). Then, $\mathsf{S}_{\mathrm{sde}}$ is the set where the coefficient of ISDE (13.14) is defined on $\mathsf{S}_{\mathrm{sde}}^{[1]}=(\mathfrak{u}^{[1,0]})^{-1}(\mathsf{S}_{\mathrm{sde}})$.
Let $(\sigma,\mathfrak{b})$ be the coefficient of ISDE (2.41). We define $\sigma^{m}$ and $\mathsf{b}^{m}$ by
$\sigma^{m}(\mathbf{x},\mathsf{s})=(\sigma({x}^{i}))_{i=1}^{m},\quad\mathsf{b}^{m}(\mathbf{x},\mathsf{s})=\Big{(}\mathfrak{b}({x}^{i},\sum_{j\neq i}^{m}\delta_{{x}^{j}}+\mathsf{s})\Big{)}_{i=1}^{m}$ (10.8)
for $\mathbf{x}=(x^{1},\ldots,x^{m})$ and $\mathsf{s}\in\mathsf{S}$. Let $\mathfrak{H}[\mathbf{a}]$ be as given by (13.20)–(13.22).
[…]
ii For each $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}$, there exists a constant $\mathsf{F}_{\mathsf{m}}$ such that
$|\sigma^{m}(\mathbf{x},\mathsf{s})-\sigma^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{m}}|\mathbf{x}-\mathbf{y}|,$ (13.27)
$|\mathsf{b}^{m}(\mathbf{x},\mathsf{s})-\mathsf{b}^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{m}}|\mathbf{x}-\mathbf{y}|$ (13.28)
for all $(\mathbf{x},\mathsf{s}),(\mathbf{y},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$ such that $(\mathbf{x},\mathsf{s})\sim_{\mathsf{m}}(\mathbf{y},\mathsf{s})$, where $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ is the set defined by (13.21) and $\sim_{\mathsf{m}}$ is the equivalence relation defined after (13.26).
[…]
Note that $\sigma\in C_{b}^{1}(\mathbb{R}^{d})$ and that $\sigma$ is independent of $\mathsf{s}$ by assumption. Hence, $\sigma^{m}$ is Lipschitz continuous, which yields (13.27).
From Proposition 5.1, we have $\mathscr{R}_{R,\mathsf{s}}^{\ell}\in C^{1}(\overline{S}_{R})$. From (5.53), we have
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big{(}-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big{)}.$ (10.9)
[…]
Theorem 10.5.
Assume ((A1))((A3)). Then, for $\mu$-a.s. $\mathsf{a}$, $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ satisfy ((IFC)).
@@
Assume ((A1))((A3)). Then, for $\mu$-a.s. $\mathsf{a}$, the following hold.
@@
ii The $m$-labeled processes $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ and $\overline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ given by $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$, respectively, are conservative $\mu_{\mathsf{a}}$-symmetric diffusions that are properly associated with identical strongly local, quasi-regular Dirichlet forms on $L^{2}(\mu_{\mathsf{a}}^{[m]})$:
$(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})=(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}}).$ (10.10)
Here, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ and $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ are given by Lemma 10.3.
@@
From Theorem 10.2 and Theorem 10.3, $\underline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ and $\overline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ are associated with the $m$-labeled Dirichlet forms $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ and $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$, respectively. Because $\underline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ and $\overline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ are indistinguishable, we obtain (10.10).
From Theorem 10.3, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ is a strongly local, conservative, and quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence, from (10.10), $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ is also a strongly local, conservative, and quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. This implies ii. ∎
@@
The method to use the uniqueness of strong solutions of ISDEs to prove the uniqueness of the extension of Dirichlet forms on interacting Brownian motions is due to Tanemura [43]. This method was followed by [15].
Proof of Theorem 2.3.
Let $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ be as in Theorem 10.6. Then, they are indistinguishable from Theorem 10.6. Hence, we write $\mathbf{X}:=\underline{\mathbf{X}}_{\mathsf{a}}=\overline{\mathbf{X}}_{\mathsf{a}}$.
@@
11. Uniqueness of the extension of Dirichlet forms: Proof of Theorem 2.4.
This section aims to prove the uniqueness of the extensions of Dirichlet forms and, using this, to establish the fully labeled process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ in Theorem 2.3 is a conservative diffusion on $(\mathbb{R}^{d})^{\mathbb{N}}$ and, thus, complete the proof of Theorem 2.4.
11.1. Uniqueness of extension of Dirichlet forms.
We now analyze the Dirichlet forms directly related to the solution of ISDE (2.41) to prove the uniqueness of extension of Dirichlet forms in Theorem 11.1.
Let $(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})$ and $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}})$ be as in Lemma 7.2. Recall that
$\mu^{[m]}=\int_{\mathsf{S}}\mu_{R,\mathsf{y}}^{[m]}\mu(d\mathsf{y}).$
Let $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ be the superposition of $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$ [4]. By definition, $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ is the closed form on $L^{2}(\mu^{[m]})$ given by
$\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}}(f,g)=\int_{\mathsf{S}}\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}}(f,g)\mu(d\mathsf{y}),$
$\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}}=\{f\in\bigcap_{\text{$\mu$-a.s.\,$\mathsf{y}$}}\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}};\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}}(f,f)<\infty,f\in L^{2}(\mu^{[m]})\}.$ (11.1)
@@
Assume ((A1))((A3)). Let $m\in\{0\}\cup\mathbb{N}$. Then,
$(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})=(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})\quad\text{ for all }R\in\mathbb{N}.$ (11.2)
@@
Let $\underline{\mathbf{X}}_{R}^{[m]}$ and $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$ be the $m$-labeled diffusions given by the Dirichlet forms $(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$ and $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$, respectively. Let $\underline{\mathbf{X}}_{R}$ and $\underline{\mathbf{X}}_{R,\mathsf{y}}$ be the fully labeled processes given by the diffusions $\underline{\mathbf{X}}_{R}^{[m]}$ and $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$, respectively. From the argument before Lemma 8.1, such processes exist. According to Lemma 8.1, both fully labeled processes are weak solutions of the pair of SDEs (2.58) and (2.59).
@@
From the representation of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})$ given by (5.5), the pair of SDEs (2.58) and (2.59) has a unique strong solution. Indeed, the coefficients given by the logarithmic derivative are smooth outside $\mathscr{C}^{[1]}=\{({x},\mathsf{s})\in\mathbb{R}^{d}\times\mathsf{S};{x}=s^{i}\text{ for some }i\}$, where $\mathsf{s}=\sum_{i}\delta_{s^{i}}$, and each particle is posed the reflecting boundary condition on $\partial S_{R}$. Hence, (2.58) has a unique strong solution until the solution hits $\mathscr{C}^{[1]}$, which never happens because of Lemma 8.1 iv. Here, Lemma 8.1 iv is clearly valid for $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[1]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[1]}})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[1]})$ similarly as for $(\mathscr{E}_{R}^{\mu^{[1]}},\underline{\mathscr{D}}_{R}^{\mu^{[1]}})$ on $L^{2}(\mu^{[1]})$. Thus, (2.58) has a unique strong solution. SDE (2.59) means that all particles outside $\overline{S}_{R}$ are frozen. Hence, it is clear that (2.59) has a trivial unique strong solution.
@@
Because $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ is the superposition of $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}})$, the collection of diffusions $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$ associated with $(\mathscr{E}_{R}^{\mu_{R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{R,\mathsf{y}}^{[m]}})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$ gives the diffusion associated with $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$. As we saw at the beginning of the proof, the diffusion $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$ gives the unique solution of (2.58) and (2.59). Hence, the diffusion $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$, $\mathsf{y}\in\mathsf{S}$, given by the superposition $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ is the unique solution of (2.58) and (2.59).

### v4

@@ 10.1. Tail decomposition of Dirichlet forms and weak solution
Let $(\sigma,\mathfrak{b})$ be the coefficient of ISDE (2.33). We define $\sigma^{m}$ and $\mathfrak{b}^{m}$ by
$\sigma^{m}(\mathbf{x},\mathsf{s})=(\sigma({x}^{i}))_{i=1}^{m},\quad\mathfrak{b}^{m}(\mathbf{x},\mathsf{s})=\Big(\mathfrak{b}({x}^{i},\sum_{j\neq i}^{m}\delta_{{x}^{j}}+\mathsf{s})\Big)_{i=1}^{m}$ (10.8)
for $\mathbf{x}=(x^{1},\ldots,x^{m})$ and $\mathsf{s}\in\mathsf{S}$.
We verify that $\sigma^{m}$ and $\mathsf{b}^{m}$ in (10.8) satisfy (13.13) and (13.14).
By assumption, $\sigma\in C_{b}^{1}(\mathbb{R}^{d})$ and that $\sigma$ is independent of $\mathsf{s}$. Hence, $\sigma^{m}$ is Lipschitz continuous, which yields (13.13).
By (5.18), we have $\mathscr{R}_{R,\mathsf{s}}^{\ell}\in C^{1}(\overline{S}_{R})$. From (2.26), we obtain
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big).$
Hence, $\mathsf{d}^{\mu}$ is Lipschitz continuous on $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$ from (13.10) and (13.11). Then from $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$, $\mathfrak{b}=\frac{1}{2}\{\mathrm{div}\mathfrak{a}+\mathsf{d}^{\mu}\mathfrak{a}\}$ is Lipschitz continuous on $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$. Hence, by (10.8), $\mathfrak{b}^{m}$ is Lipschitz continuous on $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$, which yields (13.14). ∎
Proposition 10.4.
For $\mu$-a.s. $\mathsf{a}$, $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ satisfy ((IFC)).
@@
For $\mu$-a.s. $\mathsf{a}$, the following hold:
@@
ii The $m$-labeled processes $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ and $\overline{\mathbf{X}}_{\mathsf{a}}^{[m]}$, induced by $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$, respectively, are conservative, $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusions properly associated with the same strongly local, quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$:
$(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]}).$ (10.9)
Here $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ and $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ are given by Lemma 10.4.
@@
From Propositions 10.1 and 10.2, $\underline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ and $\overline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ are associated with the $m$-labeled Dirichlet forms $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ and $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$, respectively. Because $\underline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ and $\overline{\mathbf{X}}^{[m]}_{\mathsf{a}}$ are indistinguishable, we obtain (10.9).
From Theorem 10.2, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ is a strongly local, conservative, and quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence, from (10.9), $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ is also a strongly local, conservative, and quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. This implies ii. ∎
@@
The method of using the uniqueness of strong solutions of ISDEs to prove the uniqueness of extensions of Dirichlet forms for interacting Brownian motions is due to Tanemura [43]. This approach was subsequently adopted in [17].
Proof of Theorem 2.3. Let $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ be as in Theorem 10.1. By Theorem 10.1, they are indistinguishable. Hence, we write $\mathbf{X}:=\underline{\mathbf{X}}_{\mathsf{a}}=\overline{\mathbf{X}}_{\mathsf{a}}$.
@@
For each $m\in\{0\}\cup\mathbb{N}$, the $m$-labeled process $\mathbf{X}^{[m]}$ induced by $\mathbf{X}$ is properly associated with the strongly local, quasi-regular Dirichlet form in (10.9) on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence, $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion. By Proposition 10.1, $\mathbf{X}^{[m]}$ is conservative. This completes the proof of Theorem 2.3. ∎
11. The diffusion on $(\mathbb{R}^{d})^{\mathbb{N}}$: Proof of Theorem 2.4
The aim of this section is to prove the uniqueness of extensions of Dirichlet forms and to construct the diffusion on $(\mathbb{R}^{d})^{\mathbb{N}}$. We first establish the uniqueness of Dirichlet forms and, using this result, show that the fully labeled process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ constructed in Theorem 2.3 is a conservative diffusion on $(\mathbb{R}^{d})^{\mathbb{N}}$. This completes the proof of Theorem 2.4.
Unless stated otherwise, all results in Section 11 are proved under Assumptions ((A1))((A3)).
11.1. Uniqueness of extensions of Dirichlet forms
We now analyze the Dirichlet forms directly associated with solutions of the ISDE (2.33) to prove the uniqueness of extensions of Dirichlet forms asserted in Theorem 11.1.
Let $\mu_{\mathsf{a}}$ be as in Definition 2.2. Let $\mu_{\mathsf{a}}^{[m]}$ be the $m$-Campbell measure of $\mu_{\mathsf{a}}$. Recall that
$\mu_{\mathsf{a}}^{[m]}=\int_{\mathsf{S}}\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}\mu(d\mathsf{y}).$ (11.1)
Let $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ be as in Lemma 10.3. Let $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ be the superposition of $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ with respect to $\mathsf{y}$ under the measure $\mu_{\mathsf{a}}$:
$\mathscr{E}_{\mathsf{a},R}^{[m]}(f,g)=\int_{\mathsf{S}}\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]}(f,g)\mu_{\mathsf{a}}(d\mathsf{y}),$
$\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]}=\{f\in\bigcap_{\mu_{\mathsf{a}}\text{-a.s.\,}\mathsf{y}\in[\mathsf{a}]}\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]};\mathscr{E}_{\mathsf{a},R}^{[m]}(f,f)<\infty,f\in L^{2}(\mu_{\mathsf{a}}^{[m]})\},$
where $\mathsf{a}]$ is the set in Definition 2.2. We recall that $\mu_{\mathsf{a}}([\mathsf{a}])=1$. By the general theorem on the superposition of closed forms [3], $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ is a closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Let $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ be as in Lemma 10.3.
@@ 11.1. Uniqueness of extensions of Dirichlet forms
Let $m\in\{0\}\cup\mathbb{N}$. For $\mu$-a.s. $\mathsf{a}$ and all $R\in\mathbb{N}$,
$(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})=(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]}).$ (11.2)
@@
Let $\underline{\mathbf{X}}_{\mathsf{a},R}^{[m]}$ and $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ be the $m$-labeled diffusions associated with the Dirichlet forms $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ and $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$, respectively. Let $\underline{\mathbf{X}}_{\mathsf{a},R}$ and $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}$ be the fully labeled processes associated with the diffusions $\underline{\mathbf{X}}_{\mathsf{a},R}^{[m]}$ and $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$, respectively. By the argument preceding Lemma 8.1, such processes exist and are unique.
By Lemma 8.1, both fully labeled processes $\underline{\mathbf{X}}_{\mathsf{a},R}$ and $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}$ are weak solutions of the pair of SDEs (8.1) and (8.2). From the representation of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})$ given by (5.6), the pair of SDEs (8.1) and (8.2) has a unique strong solution.
@@
Indeed, the coefficients given by the logarithmic derivative are smooth outside $\mathscr{C}^{[1]}=\{({x},\mathsf{s})\in\mathbb{R}^{d}\times\mathsf{S}\,;\,{x}=s^{i}\text{ for some }i\}$, where $\mathsf{s}=\sum_{i}\delta_{s^{i}}$, and each particle is subject to the reflecting boundary condition on $\partial S_{R}$.
Hence, (8.1) admits a unique strong solution until the solution hits $\mathscr{C}^{[1]}$, which never occurs by Lemma 8.1 iv. Here, Lemma 8.1 iv is clearly valid for $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[1]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[1]})$ on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[1]})$ as well, similarly to $(\mathscr{E}_{\mathsf{a},R}^{[1]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[1]})$ on $L^{2}(\mu_{\mathsf{a}}^{[1]})$. Thus, (8.1) admits a unique strong solution.
Equation (8.2) means that all particles outside $\overline{S}_{R}$ are frozen. Hence, (8.2) clearly admits a trivial unique strong solution.
Let $\mathrm{Cnf}_{R,\mathsf{y}}^{[m]}$ be as in the proof of Lemma 7.3. Then $\mathrm{Cnf}_{R,\mathsf{y}}^{[m]}$ is an invariant set of the diffusion $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ properly associated with $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$.
Recall that $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ is the superposition of $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ with respect to $\mathsf{y}$ under $\mu_{\mathsf{a}}$, and that $\mathrm{Cnf}_{R,\mathsf{y}}^{[m]}$ yields a partition of $(\mathbb{R}^{d})^{m}\times\mathsf{S}$. Hence the family of diffusions $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ associated with $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$ gives rise to a diffusion associated with $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Here we used (11.1).
@@
As observed at the beginning of the proof, $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ is the unique solution of (8.1) and (8.2). Hence, for each $\mathsf{y}\in\mathsf{S}$, the diffusion obtained from $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ is this unique solution.
Recall that $\underline{\mathbf{X}}_{\mathsf{a},R}$ is a weak solution of SDEs (8.1) and (8.2).